RAL Tier1 Experiments Liaison Meeting
Access Grid
RAL R89
-
-
13:30
→
13:34
Site Operations 4m
-
13:34
→
13:35
Experiment Operational Issues 1m
-
13:35
→
13:40
ATLAS Operations Report 5mSpeakers: Brij Kishor Jashal (Rutherford Appleton Laboratory), Jyoti Prakash Biswal (Rutherford Appleton Laboratory)
-
13:40
→
13:45
CMS Operations Report 5mSpeaker: Katy Ellis (Science and Technology Facilities Council STFC (GB))
The CVMFS issue of the last weeks has now been completely resolved - thanks to Alex for all his efforts. Presumably the routing issue to Brazil still exists (the solution to the CMS SAM test problem was simply to remove the Brazilian host from the list of hosts tested by RAL) but James Adams was made aware.
However, now I am seeing occasional, intermittent missing tests and failures across a whole range of different SAM tests https://cmssst.web.cern.ch/siteStatus/detail.html?site=T1_UK_RAL
These affect the CE tests only, not the storage.
Example of the CVMFS test failing:
[C] ls of /cvmfs/cms-griddata.cern.ch timed out
There are also timeouts on data access, squid and frontier tests. Does this point to some network instability?
Jobs and transfers seem to be running ok.
The folks at Imperial College decided their AAA redirectors weren't really working and they didn't want to fix them, so they have been removed from the federation today. We'll see how well the remaining redirector based at T1 copes by itself.
-
13:45
→
13:50
LHCb Operations Report 5mSpeaker: Alexander Rogovskiy (Rutherford Appleton Laboratory)
Another big staging campaign is approaching -- we are going to start staging this Friday. According to my back-of-the envelope calculation, around 2.6PB is going to be staged.
Buggy xrootd client affected ETF tests for some dcache sites last Friday, RAL was not affected.
There were some incorrect user activities which resulted in failed uploads. Also a couple of (relatively small) buggy productions were submitted by LHCb. Nothing else of note, i think. - 13:50 → 13:55
-
13:55
→
14:00
LSST Operations Report 5mSpeakers: Thomas Birkett, Timothy Noble (Science and Technology Facilities Council STFC (GB))
- Metadata Database issues regarding slowness for jobs
- Ticket in to Database team to change some settings as I no longer have access after the access list changes, and DB team have said I cannot get access
- Jobs are running OK at RAL other than the above slow database:

- Data has started to flow again from the Observatory, not yet to RAL, but that will come - IN2P3 is taking the 2nd full copy so it is the 1st port of call for any new raw data after SLAC

- Metadata Database issues regarding slowness for jobs
-
14:00
→
14:01
Tier-1 Projects 1m
-
14:05
→
14:10
XRootD Development 5mSpeakers: Alexander Rogovskiy (Rutherford Appleton Laboratory), Jyothish Thomas (STFC)
-
14:15
→
14:20
Utilizing GPUs 5mSpeakers: Dr Brij Kishor Jashal (Rutherford Appleton Laboratory), Thomas Birkett
-
14:25
→
14:45
Hardware deployment 20m
Echo SN
WN
GPUSpeaker: Jacob Ward -
14:45
→
14:46
DC 27 1mSpeaker: Katy Ellis (Science and Technology Facilities Council STFC (GB))
-
14:50
→
14:59
Summary of Operational Status and Issues 9mSpeakers: Brian Davies (Science and Technology Facilities Council STFC (GB)), Darren Moore, Thomas Birkett
-
15:00
→
15:05
Any other Business 5mSpeakers: Brian Davies (Science and Technology Facilities Council STFC (GB)), Darren Moore
-
15:15
→
15:16
AOB 1m
-
13:30
→
13:34
